### v1

# The Brain’s Bitter Lesson: Scaling Speech Decoding With Self-Supervised Learning

Abstract
The past few years have produced a series of spectacular advances in the decoding of speech from brain activity. The engine of these advances has been the acquisition of labelled data, with increasingly large datasets acquired from single subjects. However, participants exhibit anatomical and other individual differences, and datasets use varied scanners and task designs. As a result, prior work has struggled to leverage data from multiple subjects, multiple datasets, multiple tasks, and unlabelled datasets. In turn, the field has not benefited from the rapidly growing number of open neural data repositories to exploit large-scale data and deep learning. To address this, we develop an initial set of neuroscience-inspired self-supervised objectives, together with a neural architecture, for representation learning from heterogeneous and unlabelled neural recordings. Experimental results show that representations learned with these objectives generalise across subjects, datasets, and tasks, and are also learned faster than using only labelled data. In addition, we set new benchmarks for two foundational speech decoding tasks. Taken together, these methods now unlock the potential for training speech decoding models with orders of magnitude more existing data.
Figure 1: Leveraging unlabelled data for speech decoding. We use abundant unlabelled neuroimaging data to pre-train a neural network using tasks that do not require explicit labels. The tasks apply a randomly selected transformation $T$ to the data and the network predicts the transformation. We fine-tune the pre-trained network on labelled data, achieving faster generalisation owing to the strength of the representation learned in pre-training.
1 Introduction
In his Bitter Lesson, Richard Sutton argues that a major conclusion of 70 years of AI research is that general methods exploiting large-scale computation will outperform model-based approaches as the availability of compute increases [50]. In line with this, the generality of deep learning, via statistical learning from ever bigger datasets, has allowed the field to leverage computation in a way that appears to scale arbitrarily, leading to astounding advances across a diverse set of domains [29, 7, 42, 46].
In the domain of brain data, and of tasks like speech decoding, the bitter lesson has not yet fully assimilated. State-of-the-art brain-computer interfaces (BCIs) have tried to scale up labelled datasets for individual subjects, using either invasive [40, 60] or non-invasive brain recordings [52], mapping these to transcripts of attempted or imagined speech. Yet, a number of obstacles to scale remain. With few exceptions at present (e.g. [17]), speech decoding models tend not to train on data from more than one subject. Moreover, they do not combine data from multiple datasets and in general do not utilise unlabelled data, or data from diverse tasks. Thus the size of training data has been limited to how much can be acquired for a single subject, and data from other subjects, or from the growing number of public data repositories, has not been leveraged. There are many reasons for these limitations; individual brains and data from different neuroimaging scanners differ, for example. But overcoming these limitations, as has begun to happen in neighbouring sub-fields (e.g. [27]), holds the promise of training models on collective, internet-scale data.
Given the scarcity of labelled data, self-supervised learning (SSL) appears promising as an avenue for domains where such data is rare or hard to obtain [3]. In the SSL paradigm, pretext tasks pre-train a model on unlabelled data by generating implicit training labels through transformations of the input data in order to help a downstream task. We develop a set of these tasks, informed by advances in neuroscience, for learning with unlabelled brain data (Figure 1) and design an architecture for processing continuous multi-sensor neuroimaging signals. In order to scale existing non-invasive datasets, we provide a unified method that allows us to leverage data from other experiments that do not have the same labels (by treating them as unlabelled) and that come from different subjects and neuroimaging scanners. We evaluate the representations learned with our approach on heard speech datasets acquired with non-invasive magnetoencephalography (MEG), setting the baselines for speech detection and voicing classification on this data. The results not only demonstrate that scaling with unlabelled data works in speech decoding, but also shows that these representations can generalise across datasets, tasks, and even novel subjects for the first time. Our main contributions are:
  • A set of domain specific self-supervised pretext tasks for representation learning that can scale speech decoding over multiple subjects, multiple studies, and unlabelled data;
  • A neural architecture for learning these self-supervised objectives and training downstream speech decoding from brain data; and
  • A comprehensive experimental evaluation providing evidence that our approach can scale up with data and enable cross-dataset, task, and subject generalisation.
2 Related Work
Prior work in speech decoding has focused almost entirely on supervised learning with decoding models that typically do not generalise across participants or experiments. This is true both in recent state-of-the-art invasive studies [40, 39, 60, 9] and non-invasive studies [52]. These prior works have scaled up the experimental data collected within individual subjects, but are unable to leverage data from other subjects and experiments. Focusing on semantic rather than phonetic decoding, work by Tang et al. [52] is remarkable for showing an ability to generalise across labelled task data when listening to speech, imagining speech, or even watching videos. They do not, however, leverage unlabelled data and are unable to show generalisation between subjects.
Specific studies into the limitations of generalising models between subjects show that while performance decreases on average when subjects are pooled, there are exceptions. Csaky et al. [11] find that a subset of individuals perform better when evaluated with a group-level model than with individual models. Exploiting audio data in a multi-modal framework, Défossez et al. [17] show that decoding performance improves for a segment identification task as data from multiple subjects listening to connected speech are aggregated. Although they repeat the result within two MEG and two EEG datasets, Défossez et al. [17] do not show any improvements for pooling data across datasets. Moreover, they do not combine data from studies with different labels either; cf. [58, 16, 56]. Unfortunately, two of these papers [58, 16] included a bug in their evaluation code. As such, their methods may perform no better than a baseline that provides pure noise inputs to the model [28].
[…]
In the computer vision literature, there have been a plethora of methods that use self-supervised pretext tasks for representation learning [1, 15, 41, 31, 62, 21]. Until now, similar approaches have not translated to the brain decoding literature. However, prior work has used other methods to leverage unlabelled brain data. For example, Jiang et al. [27] succeeded in cross-dataset and cross-task generalisation, using a transformer with tokenised brain signals and a masked token prediction objective. Although this work combined unlabelled EEG datasets, it only achieved improvements on non-speech tasks. Wang et al. [57] used a similar approach, replacing tokens with contextualised embeddings of time-frequency input representations. They attained impressive speech detection results but with invasive neural recordings, which are comparatively rare and thus have much less potential to scale than non-invasive data.
3 Method
To process continuous neuroimaging data, we introduce a neural architecture for encoding heterogeneous brain signals into latent representations. By developing pretext tasks with the objective of learning generalisable brain representations, we leverage this architecture for self-supervised learning from unlabelled data, hoping to replicate similar successes in computer vision [21, 8].
3.1 Network Architecture
We design a two-stage neural network architecture (Figure 2). The pre-training stage uses pretext tasks to train a representation with unlabelled brain data. Then, the fine-tuning stage uses this representation to learn the downstream task by fine-tuning with labelled data.
Figure 2: Architecture overview. Inputs are projected into a shared dimension, then encoded. In pre-training, all weights are trainable except for modules in light-red, while in shallow fine-tuning, modules with light-blue borders are frozen and modules with light-red borders are unfrozen. Deep fine-tuning is identical, except the encoder is trainable. Dashed borders indicate optional components.
Normalisation. We divide recordings into windows of length $w$ seconds or $t$ samples. At train time, each batch of windows is standardised such that each sensor has zero mean and unit variance.
Backbone. The network takes as input the standardised sample windows. To combine heterogeneous datasets, which have different numbers of sensors $S$, we apply a dataset-conditional linear layer to the sensor dimension, projecting the signal into a shared space with dimension $d_{\mathrm{shared}}$. Then, to encode the signal, we construct a wave-to-wave convolutional encoder architecture, the cortex encoder, inspired by work in neural audio codecs [61, 14]. Specifically, our convolutional encoder adapts the implementation of the SEANet architecture [51] used in Défossez et al. [14]. As these codecs typically operate on mono audio signals in $\mathbb{R}^{1\times t}$, while our signals are in $\mathbb{R}^{d_{\mathrm{shared}}\times t}$, we increase the convolutional channel dimension from $1$ to match $d_{\mathrm{shared}}$ while also inflating the channel dimension of subsequent convolutions. We refer to the output dimension of embeddings from this backbone as $d_{\mathrm{backbone}}$. Thus, the backbone takes as input a window in $\mathbb{R}^{S\times t}$, and encodes this into $\tau$ embeddings (where $\tau<t$), each of dimension $d_{\mathrm{backbone}}$ (i.e. an $\mathbb{R}^{d_{\mathrm{backbone}}\times\tau}$ output).
Pre-training. Following the advice of Balestriero et al. [3, Section 3.2], we use a two-layer feedforward projector to alleviate misalignment between our pretext and downstream tasks in the representation. After the projector, linear classifiers make predictions for each of the pretext tasks.
Fine-tuning. In this stage, we introduce classifiers for the downstream tasks and train with labelled data and supervised learning (Figure 2). Depending on the experiment, we fine-tune in one of two ways. In the first, we train an MLP classifier from scratch on top of the pre-trained representation, which remains frozen. Thus, we backpropagate only through the classifier. We call this shallow fine-tuning. In the other, we fine-tune by training an MLP on top of the pre-trained representation again, but do not freeze the backbone, training end-to-end. Here, we backpropagate through the classifier and the backbone. We refer to this as deep fine-tuning. In both cases, a new trainable dataset-specific linear layer can be introduced for a novel dataset.
For speech detection, our classifier makes a prediction for each individual embedding. For voicing classification, where there is only one label for each sample window, the embeddings are flattened into a tensor in $\mathbb{R}^{d_{\mathrm{backbone}}\times\tau}$ representing the entire window. This is the input to the voicing classifier and is referred to as full epoch decoding in neuroimaging literature [12].
Subject conditioning. Just as speakers have different voices, neural responses between subjects have different characteristics. Consequently, individual variation leads to models that do not generalise well across subjects [11]. In the speech literature, models include speaker conditioning to account for these differences [20]. We take a similar approach by introducing subject conditioning, represented as a subject embedding, to our model. With a SEANet-based architecture, Zeghidour et al. [61] find that conditioning is equally as effective at the encoder bottleneck as in other stages of the model. Hence, we place ours at the cortex encoder bottleneck for simplicity. We evaluate multiple types of conditioning including subject embeddings and feature-wise linear modulation (FiLM) [44].
3.2 Pretext Tasks
Our pretext tasks are unsupervised feature learning tasks for continuous neuroimaging signals that aim to learn generalisable speech decoding features. Since different datasets use different hardware and varied numbers of sensors, we construct these tasks with labels that are agnostic to the number of sensors in the signal. This means that these tasks do not require identifying specific sensor channels.
Band prediction. In the literature, neural responses can be broadly segmented into functional frequency bands [22, 45, 36]. Delta ($\delta$) waves (0.1–4 Hz) are commonly associated with the rhythmic structure of heard speech [35], Theta ($\theta$) waves (4–8 Hz) reliably track [34] and phase-lock to the amplitude envelope of heard sentences [43], Alpha ($\alpha$) waves (8–12 Hz) relate to attentional processes and the inhibition of irrelevant information, helping to focus on relevant speech signals [49], Beta ($\beta$) waves (12–30Hz) are implicated in top-down predictive coding [5] which affects lexical processing [59], Gamma ($\gamma$) waves (30–70 Hz) occur with higher cognitive functions (e.g. memory, learning, reasoning, and planning) [19, 6], and High Gamma ($\gamma^{\mathrm{high}}$) waves ($>$70 Hz) have been linked specifically to speech detection [26] and phonemic feature classification in the STG [38] as well as phonemic feature classification in the ventral sensorimotor cortex (vSMC) [10]. As High Gamma is a relatively wide band, we have split it into two sub-bands: Lower High Gamma ($\gamma^{\mathrm{high}}_{\mathrm{lower}}$) waves (70–100 Hz) and Upper High Gamma ($\gamma^{\mathrm{high}}_{\mathrm{upper}}$) waves (100–150 Hz).
To learn representations that can distinguish between these, our band prediction task applies a band-stop filter for a randomly selected band $\omega$ to the sample $x$, passes the filtered sample $x^{\omega^{\prime}}$ through the network backbone $g$ and the corresponding predictor $f_{\mathrm{band}}$, requiring the network to predict the frequency band that was rejected. This yields the loss
$\mathcal{L}_{\mathrm{band}}=\sum_{x\in B}\mathcal{L}_{\mathrm{CE}}(f_{\mathrm{band}}(g(x^{\omega^{\prime}})),\omega),$ (1)
where $B$ is a mini-batch of samples, $\omega\in\{\delta,\theta,\alpha,\beta,\gamma,\gamma^{\mathrm{high}}_{\mathrm{lower}},\gamma^{\mathrm{high}}_{\mathrm{upper}}\}$, and $\mathcal{L}_{\mathrm{CE}}$ is the cross-entropy loss as this is a multi-class classification task.
Phase shift prediction. Phase coupling between networks of neuron populations is necessary for coordinating brain activity [18, 54]. Thus, since phase often synchronises between communicating brain areas, phase coupling between spatially distant sensors is likely to be a useful feature. Supporting this insight, recent work [27] also finds phase to be an essential component of the signal.
To learn representations that encode phase differences between brain areas, this task applies a discrete uniform random phase shift $\phi\in\{0,\frac{\pi}{8},\frac{\pi}{4},\frac{3\pi}{8},\frac{\pi}{2},\frac{5\pi}{8},\frac{3\pi}{4},\frac{7\pi}{8}\}$ to a uniformly randomly selected proportion $\rho$ of the sensors. Applying this shift to random sensors is critical since sensors are placed in different positions, capturing different regions of the brain. Uniform random selection ensures differences between any two regions of the brain are represented. The objective of this task is to predict the phase shift. This leads to a similar loss
$\mathcal{L}_{\mathrm{phase}}=\sum_{x\in B}\mathcal{L}_{\mathrm{CE}}(f_{\mathrm{phase}}(g(x^{\phi})),\phi),$ (2)
where $x^{\phi}$ describes the signal with a phase shift $\phi$ applied to a proportion of the sensors. We use a discrete number of possible phase shifts, treating it as a multi-class task rather than a regression task, to ease the difficulty of the problem as MEG scanners typically have a large number of sensors.
Amplitude scale prediction. MEG and EEG signals use an array of sensors at different spatial locations, capturing different signal sources more intensely. Representing the relative amplitude difference between sensors could be important for differentiating between neural responses originating from distinct parts of the brain. Within speech, Hamilton et al. [26] find that localised regions of the STG respond to sustained speech and speech onsets. Differentiating between neural responses from this region and others may be essential for decoding speech perception.
Thus, this pretext task focuses on learning representations that encode relative sensor amplitude differences. Similar to the phase shift task, we select a random proportion of the sensors $\rho$ and apply a discrete random amplitude scaling coefficient $A\in[-2,2]$, discretised into 16 scaling factors, to the signal. The objective is to predict the scaling factor, leading to the loss
$\mathcal{L}_{\mathrm{amplitude}}=\sum_{x\in B}\mathcal{L}_{\mathrm{CE}}(f_{\mathrm{amplitude}}(g(x^{A})),A),$ (3)
where $x^{A}$ is the signal scaled with $A$.
These pretext tasks capture complementary time- and frequency-domain properties of the signal. Hence, during pre-training, we combine them, creating an augmented version of the input for every pretext task by applying the matching transformation. We feed the augmented inputs through the network backbone and apply the corresponding classifier to predict the transformation, summing the weighted losses such that our final pre-training loss is given by
$\mathcal{L}_{\mathrm{SSL}}=w_{1}\mathcal{L}_{\mathrm{band}}+w_{2}\mathcal{L}_{\mathrm{phase}}+w_{3}\mathcal{L}_{\mathrm{amplitude}},$ (4)
where $w_{i}$ is a constant coefficient for each self-supervised loss.
4 Experiments
In this section, we evaluate the representations learned with our pretext tasks by measuring their ability to scale downstream performance with unlabelled data. This includes understanding how well they can generalise across datasets, subjects, and tasks. We focus our evaluation on MEG data as the signal is rich, with better spatial resolution than EEG [32] and faster sampling rates than fMRI [25].
4.1 Experimental setup
Datasets. Unless specified otherwise, our experiments use Cam-CAN [48, 53] as an unlabelled representation learning dataset for pre-training. This is a study containing 641 subjects with resting and sensorimotor tasks, totalling approximately 160 hours of MEG recordings. For our downstream tasks, we use two labelled heard speech MEG datasets where participants listen to short stories or audiobooks. Armeni et al. [2] contains 3 subjects who listen to 10 hours of recordings each (30 hours total) while Gwilliams et al. [24] has 27 subjects, each recorded for 2 hours (54 hours total). Overall, we utilise over 200 hours of data. To the best of our knowledge, this is the largest volume of MEG data ever used for speech decoding.
Preprocessing. Each recording is in $\mathbb{R}^{S\times T}$ where $S$ is the number of sensors and $T$ is the number of time points sampled by the scanner. To eliminate high-frequency muscle movement artifacts, we apply a low-pass filter at 125Hz as well as a high-pass filter at 0.5Hz to remove slow-drift artifacts. Since the datasets were recorded in Europe, where the electric grid frequency is 50Hz, we apply a notch filter at multiples of 50Hz to account for line noise. Next, we downsample the signal to 250Hz, avoiding aliasing at frequencies up to our low-pass filter threshold. Finally, we detect bad sensor channels, those with significant noise and artifacts, using a variance threshold and replace them by interpolating the spatially nearest sensors.
Downstream tasks. We evaluate our methods with two fundamental speech decoding tasks of increasing difficulty. The first, speech detection, determines whether speech occurs in the auditory stimulus using the neural response. The second task is voicing classification. Given data aligned at the occurrence of a phoneme, the task is to recognise whether the phoneme is voiced or voiceless, where voicing is a binary phonetic feature that categorises whether a speech sound is associated with vocal cord vibration. We select these tasks as they are simpler than phoneme recognition, but are foundational because they must be solved to decode speech accurately into natural language.
Training. We pre-train all models to completion and then shallow or deep fine-tune on labelled data for each task, using this to analyse the generalisability of the pre-trained representation. Appendix B provides complete training details for all experiments.
4.2 Learning Generalisable Representations Using Pretext Tasks
To understand whether the pretext tasks produce useful representations, we investigate how they generalise independently with shallow fine-tuning (i.e. with the representation frozen). In Table 1, we show that all of our pretext tasks lead to better than chance accuracy on speech detection, while for voicing the band prediction task does best. In all cases, the tasks either match or beat a randomly initialised and shallow fine-tuned baseline. The results are statistically significant whenever we beat the baselines and provide initial evidence that all of our tasks are useful for learning representations that generalise downstream to speech decoding.
Pretext task Fine-tuned balanced accuracy
Speech $t$ $p$ Voicing $t$ $p$
None (random backbone) $0.500\pm 0.000$- -- $0.500\pm 0.000$- --
Phase($\rho=0.5$) $0.538\pm 0.002$ $18.5$ $0.001$ $0.500\pm 0.000$- --
Amplitude($\rho=0.2$) $\mathbf{0.555}\pm 0.002$ $23.3$ $0.001$ $0.500\pm 0.000$- --
Band $0.551\pm 0.005$ $10.0$ $0.005$ $\mathbf{0.503}\pm 0.001$ $4.1$ $0.03$
Chance-level $0.5$- -- $0.5$- --
Table 1: Pre-training with pretext tasks produces generalisable representations. We provide the test accuracy at the point of best validation accuracy (early stopping). In the none baseline, the network backbone is randomly initialised before fine-tuning (i.e. there is no pre-training). The uncertainty is the standard error of the mean over three random seeds. We quote the $t$-score and $p$-value for each result from a one-sample one-sided $t$-test where the population mean is chance-level accuracy. Models are trained on all subjects in Armeni et al. [2] and evaluated on held-out sessions.
The strength of band prediction in voicing classification may be due to discriminating between theta waves and other frequency bands as theta waves are associated with phonetic feature processing. The choice of the proportion of sensors to apply transformations to, $\rho=0.5$ for phase shift prediction and $\rho=0.2$ for amplitude prediction, were determined through a hyperparameter search (Appendix C). We conjecture that a smaller $\rho$ is optimal for amplitude scale prediction since this leads to representations that are especially strong at discriminating amplitude differences among small groups of sensors. Perhaps this makes it easier to distinguish between neural responses from distinct parts of the brain such as the STG, which is associated with speech onset [26]. In contrast, a larger $\rho$ for phase shift prediction may be best because phase is descriptive of neural synchrony which is distributed across the brain. Unlike localised responses, a large proportion of the sensors can detect this feature.
4.3 Scaling Speech Decoding With Unlabelled Data
Previously, we pre-trained and fine-tuned using data from Armeni et al. [2]. This setting was useful to understand whether our pretext tasks benefited the representation but did not demonstrate if they were effective in a real scarce data setting, where representations must generalise across datasets in order to scale. Furthermore, we did not investigate the scaling behaviour with downstream tasks. Here, we pre-train with Cam-CAN [48, 53], which does not have any speech decoding labels, and fine-tune for ten epochs on Armeni et al. [2] and Gwilliams et al. [24], showcasing transfer with minimal fine-tuning. This simulates the scenario in which models need to generalise quickly with very little calibration data from new patients in order to be deployed as a speech BCI. We use shallow fine-tuning for speech detection and deep fine-tuning for voicing. As evidenced by Table 1, we require the latter because voicing classification is a more difficult task, necessitating end-to-end training with more parameters to observe generalisation.
Figure 3: Scaling unlabelled data improves downstream generalisation. We pre-train the model on increasing amounts of unlabelled data and fine-tune on voicing and speech detection. No pre-training shows best test accuracy for a randomly initialised network fine-tuned for the same number of epochs (i.e. without pre-training). We show only one no pre-training line as the result is the same for both datasets. The shaded area shows the standard error of the mean across three seeds.
Figure 3 shows balanced accuracy as we increase the amount of unlabelled data in pre-training up to approximately 160 hours. For both tasks, pre-training outperforms the no pre-training baseline, which fails to generalise at all, showing that any unlabelled data is sufficient to pre-train a useful representation. In both tasks, there is a clear improvement in accuracy as the amount of unlabelled data increases. Thus, adding unlabelled data improves generalisation and scales performance. With speech detection, the improvement in accuracy is more significant for Armeni et al. [2]. Noting that there is more within-subject data in this dataset than Gwilliams et al. [24], speech detection may be more sensitive to subject variation than voicing classification.
We scaled up the pre-training dataset by increasing the number of subjects. Since this led to consistent and almost monotonic improvements in downstream accuracy, our self-supervised method is an exception to the common consensus that pooling subjects worsens generalisation. Moreover, as we pre-trained our model with a different dataset to those we fine-tuned on, our representation shows cross-dataset generalisation. This is particularly notable as the Armeni et al. [2], Gwilliams et al. [24], and pre-training datasets all use different scanners entirely. Performing well across these datasets indicates that our pretext tasks successfully generate representations that are generalisable across heterogeneous scanners. Beyond hardware, our results also show improvements on distinct tasks, thus achieving cross-task generalisation. Finally, our pre-training dataset contained no language data whatsoever yet still improved downstream accuracy on language tasks, showing that remarkably any unlabelled data can be used to scale.
Whilst we have shown that our representations generalise faster than without pre-training, do they outperform supervised baselines when trained to saturation? While voicing requires deep fine-tuning to see a substantial learning signal, the speech detection task, with which we use shallow fine-tuning, can be compared directly to similarly parameterised supervised classifiers. Table 2 provides this comparison. A representation pre-trained with 160 hours of unlabelled data significantly outperforms chance, a randomly initialised backbone, and supervised training on both downstream datasets. Here, the supervised classifier has significantly more parameters than the models that probe the network backbone because the input dimension is much larger without an encoder. Even with this bias towards the supervised classifier, the self-supervised case performs better. Thus, it is apparent that pre-training with unlabelled data significantly improves downstream generalisation.
Model Speech detection balanced accuracy
Armeni $t$ $p$ Gwilliams $t$ $p$
Supervised
Linear $0.5222\pm 0.0015$ $12$ $0.003$ $0.5006\pm 0.0001$ $5$ $0.02$
Self-supervised + fine-tuning
Random backbone
   MLP (two layers) $0.5000\pm 0.0000$- -- $0.5000\pm 0.0000$- --
Pre-trained backbone
   Linear $0.5853\pm 0.0017$ $52$ $0.0003$ $0.5011\pm 0.0001$ $9$ $0.006$
   MLP (two layers) $\mathbf{0.6101}\pm 0.0017$ $51$ $0.0002$ $\mathbf{0.5230}\pm 0.0002$ $113$ $0.00004$
Chance-level $0.5$- -- $0.5$- --
Table 2: Self-supervised pre-training outperforms supervised classifiers. We use shallow fine-tuning on top of a randomly initialised and pre-trained backbone, comparing these to a supervised classifier trained to completion. We quote test accuracy at the point of best validation accuracy (early stopping) and show uncertainty as the standard error of the mean from three seeds. We calculate the $t$-score and $p$-value from one-sample one-sided $t$-tests where the population mean is chance.
4.4 Generic Self-Supervised Representations Improve Speech Decoding
Figure 4: Cross-task semi-supervised learning does not improve generalisation. We pre-train with increasing amounts of labelled data from Armeni et al. [2], using speech labels when evaluating voicing and voicing labels for speech. We also use a constant 38 hours of unlabelled data in pre-training. No pre-training shows test accuracy for a randomly initialised network fine-tuned for the same number of epochs (i.e. without pre-training). SSL is when pre-training with only the unlabelled data. The shaded area shows the standard error of the mean across three seeds.
Our self-supervised tasks led to clearly generalisable and scalable representations by providing an inductive bias towards speech decoding. However, could these representations be made more powerful by utilising labels from other speech decoding tasks to improve this bias? This moves us into the semi-supervised learning regime, where we use a mix of labelled and unlabelled data.
With the same experimental setup as Figure 3, we add speech detection or voicing as an additional loss during pre-training. Figure 4 shows that no amount of additional labelled data improves downstream accuracy, with all data points remaining near the range of the standard error of the mean of the SSL baseline for both tasks. There also does not appear to be any indication of scaling within the range of labelled data available to us.
This outcome indicates that our self-supervised losses are the main contributor to downstream accuracy. The MEG signal appears rich, encapsulating lots of information, with the speech component being a subset distributed across the representation space. Learning with specific labels produces representations that are too specialised to improve on other speech decoding tasks. This does not preclude there being some—or a combination of—labelled tasks that provide more general inductive biases for speech. However, the overall point remains salient. The generic property is lost with specific labels while it is well-preserved with SSL, providing another reminder of the bitter lesson.
4.5 Generalising Zero-Shot To Novel Subjects
Brain data is variable across participants, leading to difficulty transferring models to novel subjects [11]. Whilst we have shown generalisation across subjects with our method, here, we investigate how well we can generalise to novel subjects—an even more difficult challenge. We pre-train then fine-tune our models on a subset of Gwilliams et al. [24], holding out three subjects for evaluation.
We conduct these experiments without conditioning, with subject embeddings, and with FiLM to see if they have any influence on subject generalisation. We represent embeddings as learnable vectors, unique to each subject, that are concatenated to the backbone representation. For FiLM, we use the same embeddings as inputs to the functions that modulate the backbone representation.
Subject Subject voicing balanced accuracy
conditioning Seen $t$ $p$ Unseen (zero-shot) $t$ $p$
None $0.5118\pm 0.001$ $8.4$ $0.007$ $0.5057\pm 0.001$ $4.5$ $0.02$
Embedding $0.5168\pm 0.0006$ $27.9$ $0.0006$ $0.5052\pm 0.0007$ $7.5$ $0.009$
FiLM [44] $\mathbf{0.5197}\pm 0.001$ $18.6$ $0.001$ $\mathbf{0.5058}\pm 0.002$ $3.3$ $0.04$
Chance-level $0.5$- -- $0.5$- --
Table 3: Self-supervised representations generalise zero-shot to novel subjects. We quote test accuracy at the point of best validation accuracy (early stopping). Uncertainty is the standard error of the mean and we show the $t$-score and $p$-value from a one-sample one-sided $t$-test where the population mean is chance-level accuracy.
Table 3 reveals that in-distribution subject generalisation is helped by using subject conditioning methods. Both subject embeddings and FiLM conditioning outperform no conditioning, with FiLM being best. However, for novel subject generalisation, using a subject conditioning method makes little difference. Regardless of conditioning, we can generalise beyond chance-level, but suffer a reduction in accuracy compared to in-distribution subjects. Conditioning may be negligible in the unseen case as the embeddings are random rather than learned for the held-out subjects.
4.6 Limitations
Although our results are significant in demonstrating a viable path forward to scale up speech BCIs, there remain a number of limitations to the present work. We focused here on two downstream tasks: speech detection and voice classification. Ultimately, we would like to expand this work to predict full transcripts from brain recordings (i.e. brain-to-text). This has been achieved with surgical data [40, 60] but not yet convincingly with non-invasive methods like MEG or EEG [28]. Speech detection has played an important role in the development of full brain-to-text in a surgical context [40] and we hope may play a similar role for non-invasive methods. Prior work has further used voice classification as a stand in for phoneme classification [23], and we have been able to improve on these results here. In future work, we would like to expand this to all English phonemes. Secondly, while we have been able to demonstrate the utility of a few pretext tasks here, we do not claim to have exhausted the full set of useful tasks. Rather, we conjecture that more useful pretext tasks remain to be found and believe a useful avenue of research will be into other input representations for brain recordings. For example, this paper did not make use of spatial features. Another limitation is our emphasis on heard speech over other types of speech, such as attempted or imagined speech. We hypothesise that the same methods presented here will generalise to these other varieties of speech, though this has yet to be shown. But, perhaps the biggest limitation of the present work is that, while it surpasses the amount of data used in other studies, it remains to be seen how much speech decoding tasks can be improved by scaling up the number of datasets used in training. In sharing this work now, we believe that the current proof of concept will be sufficiently impactful to the field as we continue to actively scale up the datasets that we can leverage.
5 Conclusion
Ultimately, solving speech decoding could transform the lives of patients with severe communication difficulties. Yet, this promise has not materialised because the field has been blocked by its inability to scale up data to leverage deep learning. Prior methods have been unable to aggregate data across different datasets, labels, or subjects to scale up because of heterogeneity in recording hardware, experiment design, and participants. A handful of studies have shown weak signals towards alleviating these issues. But until now, no one has developed a general solution. We provided a unified method that leverages unlabelled data using generic pretext tasks that shows that all of these problems can be solved. We verified this with experiments showing that our method not only scales with heterogeneous data but even generalises across datasets, subjects, and tasks. Our method unlocks the potential of the bitter lesson, providing a general method to exploit more computation by using more data. We implore the research community to employ the vast quantities of data and compute available to realise this potential. If scale is all you need in speech decoding, then the bitter lesson may not be so bitter.
Acknowledgments and Disclosure of Funding
We would like to thank Botos Csaba for many insightful discussions and creative ideas which helped shaped the direction of this work. Thanks also to Minqi Jiang for an encouraging conversation on unsupervised representation learning, Mats W.J. van Es for technical assistance with the OSL library, Brian Liu for technical contributions which did not reach the final paper, and Miran Özdogan for reviewing a draft of this work. The authors would like to acknowledge the use of the University of Oxford Advanced Research Computing (ARC) facility in carrying out this work. http://dx.doi.org/10.5281/zenodo.22558.
[…]
Appendix A Supervised Heard Speech Decoding
While this work is primarily about scaling speech decoding, we also set the first baselines for speech detection and voicing classification from MEG with connected auditory speech stimuli. Therefore, in this section we provide fully supervised results where our architecture is trained to completion using all available labelled data in the corresponding dataset (except the validation and test splits) and compare this to a linear classifier baseline. These results are not comparable with our self-supervised models because the entire architecture is trained end-to-end. In our SSL work with shallow fine-tuning, only the final classifier is trained with labels.
Experiment Balanced accuracy
Speech detection Voicing
Armeni (linear) $0.5222\pm 0.0015$ $0.5207\pm 0.0007$
Armeni (ours) $\mathbf{0.7253}\pm 0.0005$ $\mathbf{0.5282}\pm 0.0006$
Gwilliams (linear) $0.5006\pm 0.0001$ $0.5166\pm 0.0002$
Gwilliams (ours) $\mathbf{0.5946}\pm 0.0016$ $\mathbf{0.5211}\pm 0.0005$
Table 4: Supervised baselines. Uncertainty is standard error of the mean over three random seeds.
The results in Table 4 show that we can achieve strong speech detection results on Armeni et al. [2] and Gwilliams et al. [24] with our architecture. The difference between the two may be due to a combination of different scanners, numbers of sensors (208 vs 269 magnetometers), varied difficulty in the audio stimulus, and more within-subject data in Armeni et al. [2]. We also achieve good results on voicing classification for both datasets, easily outperforming the linear models.
One potential way to improve on both supervised and self-supervised approaches is by using a temporally aware model such as an LSTM or an attentional model such as a Transformer as the classifier. We do not investigate this within this paper as it is out of scope. Rather, we choose to use simple classifiers as the purpose of this paper is to demonstrate the potential of scaling.
Appendix B Experiment Details
To generate the data in Table 1, we pre-train our representations with non-overlapping sample windows from all subjects and sessions 001 to 009 of Armeni et al. [2] for 200 epochs with a $0.8:0.1:0.1$ train/val/test split. We shallow fine-tune with labels for 30 epochs using all subjects, but only session 010, which was held-out during pre-training.
For the data in Figure 3, we use the same number of epochs and the same ratios for data splitting. We pre-train only with Shafto et al. [48], Taylor et al. [53]. We adjust the amount of unlabelled data used by increasing the number of subjects in the sequence 1, 2, 4, 8, 17, 36, 74, 152, 312, and 641, successively randomly selecting more subjects to include. Each seed uses a different set of subjects to reduce negative effects from outlier subjects. We fine-tune with all data in Armeni et al. [2] and Gwilliams et al. [24] with the same train/val/test ratios as before.
For the results shown in Figure 4, we use the same setup as in Figure 3. In addition to the unlabelled data, we add labelled data from Armeni et al. [2], increasing the number of sessions used in the sequence 1, 3, 6, 9, 12, and 27. For this experiment, we calculate the number of labelled data hours using an estimate of 1 hour per session per subject.
Finally, for the results in Table 3, we pre-train and deep fine-tune to saturation with subjects 01-07 of Gwilliams et al. [24] and test zero-shot generalisation using held-out subjects 08-10. All other experimental details are the same as the previous two experiments.
In all experiments, we use three randomly selected seeds for each pre-training and corresponding fine-tuning run. For speech detection, since our encoder reduces the temporal dimension from 125 samples (the number of samples in a 0.5 second window with a sample rate of 250Hz) down to 5 embeddings, we downsample our speech detection labels to match using PyTorch’s torch.nn.functional.interpolate. Therefore, each speech detection label represents a 0.1 second period of time.
Appendix C Hyperparameters
We conduct a search over hyperparameters of interest to optimise our self-supervised objectives and neural architecture. In all experiments in this section, we measure downstream performance (using shallow fine-tuning) with best validation accuracy for speech detection unless explicitly specified, training and evaluating only on data from subject 001 of Armeni et al. [2] to minimise training time, over three random seeds. All uncertainties in this section are standard deviations and not standard error of the mean as in the main body.
Table 5 shows the results of a search over the proportion $\rho$ of sensors to apply the phase and amplitude pretext task transformations to. We only search up to $\rho=0.5$ as anything above this is equivalent to applying the inverse transformation to a proportion $1-\rho$ of the sensors. We find that for phase shift prediction, $\rho=0.5$ is optimal, while for amplitude scale prediction $\rho=0.2$ is best.
Task $\rho$ Balanced accuracy
Phase shift prediction 0.1 $0.5452\pm 0.0048$
0.2 $0.5566\pm 0.0014$
0.3 $0.5578\pm 0.0021$
0.4 $0.5530\pm 0.0014$
0.5 $\mathbf{0.5655}\pm 0.0026$
Amplitude scale prediction 0.1 $0.5440\pm 0.0012$
0.2 $\mathbf{0.5796}\pm 0.0050$
0.3 $0.5625\pm 0.0063$
0.4 $0.5684\pm 0.0016$
0.5 $0.5650\pm 0.0026$
Table 5: Best $p$ for phase and amplitude pretext tasks.
In our architecture, we define $d_{\mathrm{backbone}}$ to be the bottleneck dimension of the cortex encoder (i.e. the embedding dimension) as well as the size of $d_{\mathrm{shared}}$. Balestriero et al. [3], using the results from work by Bordes et al. [4], suggest that increasing the backbone output dimension in self-supervised learning leads to better downstream transfer. We investigate this for our architecture by varying the encoder dimension in Table 6. Our results are consistent with their findings. Balanced accuracy is highest at the largest encoder dimension we evaluated. Encoder dimensions larger than $2048$ led to unstable training and were not feasible given the compute resources available to us.
$d_{\mathrm{backbone}}$ Balanced accuracy
$128$ $0.5775\pm 0.0037$
$300$ $0.5715\pm 0.0012$
$512$ $0.5835\pm 0.0016$
$1024$ $0.5800\pm 0.0050$
$2048$ $\mathbf{0.5987}\pm 0.0024$
Table 6: Optimal encoder dimension.
In Table 7, we show the effect of varying the length $w$, i.e. the window size of each data sample. Increasing the window to five seconds led to unstable losses in pre-training so we did not evaluate above a window length of two seconds. This is likely to be because increasing the window size reduces the number of training samples when only non-overlapping samples are used. For speech detection, there appears to be a U-shaped distribution, where very short and very long windows lead to better downstream accuracy. For voicing classification, the window lengths we evaluated seem to make little difference. However, windows smaller than 0.1 seconds may lead to significantly lower accuracy as most phonemes will occur for at least this length of time in the neural response [23].
Window length (s) Balanced accuracy
Speech detection Voicing classification
$0.2$ $0.5919\pm 0.0014$ $0.5033\pm 0.0014$
$0.4$ $0.5804\pm 0.0036$ $\mathbf{0.5077}\pm 0.0027$
$0.5$ $0.5715\pm 0.0012$ $0.5055\pm 0.0023$
$1.0$ $0.5790\pm 0.0012$ $0.5062\pm 0.0020$
$2.0$ $\mathbf{0.6116}\pm 0.0019$ $0.5065\pm 0.0011$
Table 7: Ideal sample window length.
We also evaluated several different configurations of the SEANet architecture for the cortex encoder, converging on convolutional blocks with channel dimensions $(512,512,512,512)$.
While these ablations indicate a theoretically ideal architectural configuration, in practice, we altered our final experimental architecture due to instabilities during training when data was scaled up. Our final architecture hyperparameters achieve a balance between the best values from our hyperparameter search and stable training. These values are detailed in Table 8.
Hyperparameter Value
Window length (s) $0.5$
$\rho$ (phase) $0.5$
$\rho$ (amplitude) $0.2$
$\{w_{1},w_{2},w_{3}\}$$\{1.0,1.0,1.0\}$
$d_{\mathrm{shared}}$ $512$
$d_{\mathrm{backbone}}$ $512$
SEANet convolution channels $(512,512,512,512)$
SEANet downsampling ratios $(5,5,1)$
FiLM conditioning dimension $16$
Subject embedding dimension $16$
Pre-training epochs $200$
Optimizer AdamW [33]
Learning rate $0.000066$
Train ratio $0.8$
Validation ratio $0.1$
Test ratio $0.1$
Table 8: Experimental hyperparameters.
Appendix D Compute Resources
All experiments were run on individual NVIDIA V100 GPUs with up to 32GiB of GPU memory on a system with up to 128GiB of RAM. Each pre-training run with the maximum amount of pre-training data took approximately 150 hours (6.25 days). Fine-tuning following pre-training took up to another 12 hours. We estimate that we used approximately 2000 hours of compute for the final experimental runs, including hyperparameter searches. In total, over the course of developing this work from idea to final paper, we used around 10,000 hours of GPU compute.
Appendix E Licences For Datasets And Code
The Armeni et al. [2] dataset is distributed under CC-BY-4.0 while the Gwilliams et al. [24] dataset is distributed under the CC0 1.0 Universal licence. The Schoffelen et al. [47] dataset is distributed with a RU-DI-HD-1.0 licence from the Donders institute. The licence for the Cam-CAN [48, 53] dataset is unknown. The SEANet code adapted from Défossez et al. [14] is distributed under the MIT licence, and the OSL library, which we use for preprocessing, is under the BSD-3-Clause licence.
Appendix F Broader Impacts
Decoding speech from non-invasive brain recordings is likely to bring about significant positive societal impacts. Research in this field could enable paralysed patients to communicate freely and materially assist those with minor communication difficulty (e.g. stammering). As the technology matures, it could also enable new ways of communicating with others and interacting with devices without the risks of invasive surgical implants. Nevertheless, the maturity of this technology could also present potential negative societal impacts. For one, reading inner speech creates new concerns over data controls as this information is likely to be highly sensitive and personal to individuals. Given access to this technology, there is also the risk that bad actors could extract sensitive information from target individuals without consent. Moreover, there are possible long horizon effects associated with speech decoding research. Broad adoption of this technology could lead to the gradual erosion of privacy over inner speech within society. In addition, asymmetric effects, where some individuals or organisations can read inner speech but others are unable to, could worsen societal inequality. Within the scope of this paper, we mitigate risks associated with inner speech by focusing on decoding heard speech where there is low potential for abuse. Nonetheless, we acknowledge that this is still a stepping stone towards solving inner speech decoding.